2. In a random sample of 432 employees at non-profit organizations, 387 said they were highly satisfied with their job. At
Answer #1

EXCEL (Mega Stat) output: Hypothesis test for two independent proportions - p1 0.8958 387/432 387. 432 P2 Po 0.7405 0.8104 p(c). = 0.8958 (d). xz = 391 (e). n = 528 (1). 2 =0.7405 SE, = SE + SE = V(0.0147)* +(0.0191) = 0.0005809 = 0.0254 (g). Ô1 - 2